Calculate the linear momentum, kinetic energy and total energy (in J and MeV) of a particle travelling with speed 0.90c given that its rest mass 1.68×10⁻²⁷ kg and 1MeV = 1.6×10⁻¹³ J
1
Expert's answer
2021-10-31T18:12:14-0400

Given:

"v=0.90c"

"m=1.68*10^{-27}\:\rm kg"


The linear momentum

"p=\frac{mv}{\sqrt{1-\frac{v^2}{c^2}}}"

"p=\frac{1.68*10^{-27}\:\rm kg*0.90*3*10^8\: m/s}{\sqrt{1-0.90^2}}=1.04*10^{-18}\:\rm kg\cdot m/s"

The total energy

"E=\frac{mc^2}{\sqrt{1-\frac{v^2}{c^2}}}"

"E=\frac{1.68*10^{-27}\:\rm kg*(3*10^8\: m/s)^2}{\sqrt{1-0.90^2}}\\
=3.47*10^{-10}\:\rm J=2168\: MeV"